OpenStudy (anonymous):

An equilateral triangle has an area of 36 square root 3 inches squared.what is the length of one side

OpenStudy (anonymous):

12 inches

OpenStudy (ash2326):

area of equilateral traiangle = root (3 ) *a ^2/4 we have area= 36 root (3) equating the two 36 root 3= root 3 * a^2/4 144= a^2 a=12 inches so each side is of length 12 inches